PMP– Project Management

You have to evaluate an Infrastructure Investment project in the P.R. of China. To make a solid recommendation you have to calculate the NPV, Payback Period, and the IRR as we discussed in class.

You have two scenarios you need to evaluate before making your final recommendation.

Alternative 1

The first option includes an initial investment of 19,500,000 and an annual maintenance fee of 500,000 at the end of each year. The rate of return i = 0.1 (10%). The revenue is shown below.

Calendar year    Year Cash Outflow Ao Cash inflow Ft  

2011 0. 19,500,000   

2012 1. 500,000. 3,950,000

2013. 2. 500,000. 6,010,000

2014. 3. 500,000. 7,925,000

2015. 4. 500,000. 10,300,000

2016. 5. 500,000. 12,000,000

Alternative 2

The second option includes an initial investment of 19,500,000 as well, but does not require a maintenance fee at the end of each year. The rate of return i = 0.1 (10%). However, the revenue is fixed at 6,755,000 for each of the five years.

Calendar year    Year Cash Outflow Ao Cash inflow Ft  

2011 0. 19,500,000   

2012 1.     6,755,000

2013. 2. 6,755,000

2014. 3. 6,755,000

2015. 4. 6,755,000

2016. 5. 6,755,000

           

  1. a) After you calculate the NPV, Payback Period, and the IRR for each option, make a recommendation regarding which option you would recommend and why.

  2. b) Your chief economist is telling you that an inflation rate of 5% should be expected. What is the impact on your analysis?

  3. c) Perform a sensitivity analysis to test your conclusions. Would that change your opinion or not?
Answer #1

A) NOTE - WILL BE EXCLUDING THREE ZERO'S FOR CALCULATION

NPV AND IRR (IN THE IMAGE UPLOADED)

025 B с D E F G H I K L o M N P 0 R A OPTION 1 J OPTION 1 1 NPV IRR YR 0 1 2 3 4 5 nmin 00 -19500 YR NET CASHFLOW D RATE (10%

IRR FORMULA (UPLOADED IMAGE)

usl chey yordat ith maximisation FORMULA FOR IRR IRR LR + (HK - LR) NOV (LR) [NPV (ER) - NPV (HR LR - Lower Discount Rate HR

PAYBACK PERIOD

OPTION 1

YR NET CASHFLOW BALANCE
0 -19500 -19500
1 3450 -16050 ie; (19500-3450)
2 5510 -10540
3 7425 -3115
4 9800 0

SO IT WILL BE BETWEEN YEAR 3 AND 4;

= 3+(3115/9800)

=3.318 YEARS OR 3 YEARS AND 4 MONTHS .

OPTION 2

YR NET CASHFLOW BALANCE
0 -19500 -19500
1 6755 -12745
2 6755 -5990
3 6755

IT WILL BE BETWEEN YEAR 2 AND 3

= 2+(5990/6755)

=2.886 YEARS OR 2 YEARS AND 11 MONTHS .

RESULTS SHOWS

OPTION 1 IS HAVING NPV OF 7602000, IRR OF 21.69 % AND PAYBACK PERIOD OF 3.32 YRS.

OPTION 2 IS HAVING NPV OF 6107000, IRR OF 21.7% AND PAYBACK PERIOD OF 2.89 YRS.

BOTH OPTION SHOWS POSITIVE NPV ,SIMILAR IRR , BUT OPTION 2 IS HAVING A SHORTER PAYBACK PERIOD WHICH MEANS ,WE CAN COVER INVESTMENT COST MORE FASTER THAN OPTION 1.

C) SENSITIVITY ANALYSIS

ON SALES REVENUE

OPTION 1

YR D FACTOR (10%) SALES PV
1 0.909 3950 3590
2 0.826 6010 4964
3 0.751 7925 5952
4 0.683 10300 7035
5 0.621 12000 7452
TOTAL 28993

SENSITIVITY = NPV / PV OF SALES REVENUE

= 7602/28993

=26.22%

IF CASH INFLOW IS DECREASED BY BY 26.22% ; NPV BECOMES ZERO

OPTION 2

SINCE SALES REVENUE IS SAME FOR EVERY YEAR WE CAN FIND PV OF SALES REVENUE USING ANNUITY FACTOR ( 10 % FOR 5 YEARS IN THE ANNUITY TABLE)

PV OF SALES REVENUE = 6755 * 3.791 = 25608

SENSITIVITY = 6107/25608 =23.8%

IF CASH INFLOW IS DECREASED BY BY 23.8% ; NPV BECOMES ZERO

SO OPTION 2 IS MORE SENSITIVE OR CRITICAL.

IF YOU HAVE ANY DOUBTS REGARDING THE ANSWER OR IF YOU NEED ANY CLARIFICATION

PLEASE REACHOUT ME IN THE COMMENT AND IF YOU LIKE THE ANSWER PLEASE GIVE ME A LIKE, IT WILL BE GRATEFUL.

THANK YOU